OpenStudy (anonymous):

solve the following system of quadratic and linear equations y=-x^2+4x-3 and x+y=1

OpenStudy (mathmale):

What have you thought of doing, so far? Two possible ways of doing this problem: Solve x+y=1 for either x or y. If you solve for x, then substitute the resulting expression in y into the quadratic equation. If you solve for y , then subst. the res. expr. in x into the quad. eq'n.

OpenStudy (campbell_st):

rewrite the linear equation as y = 1 - x and equate them \[1 - x = x^2 +4x + 3\]

OpenStudy (campbell_st):

simplify and solve for x

OpenStudy (anonymous):

this isnt for me its for my girl

OpenStudy (anonymous):

and who should i medal

OpenStudy (mathmale):

Evictu: Could you please help "your girl" to open her own OpenStudy account and to ask her own questions here? Your obtaining the answer and giving it to her won't help her much in the long run. Nobody's earned a medal yet.